OpenStudy (anonymous):

Jordan is a manager of a car dealership. He has 3 professional car washers to clean the entire lot of cars, Jennifer, Arianna, and Matthew. Jennifer can wash all the cars in 14 hours. Arianna can wash all the cars in 11 hours. Matthew is new to the car dealership, so no one knows how long it will take him. Jordan assigns all of them to wash the cars together. Explain to Jordan how this task can tell him how long it would take Matthew to complete the task if he worked by himself. Use complete sentences.

OpenStudy (anonymous):

I know I start with 1/11 + 1/14 + 1/m = 1/x

OpenStudy (anonymous):

@Directrix @ganeshie8

Directrix (directrix):

@hellokitty01 What does x represent in the equation you wrote?

OpenStudy (anonymous):

the time it takes for all of them to wash the cars

OpenStudy (anonymous):

This is what I got so far 25mx +154x = 154m

OpenStudy (anonymous):

And i factored x out to get x(25m + 154) = 154m

Directrix (directrix):

Take a look at the discussion here and see if you follow it: http://openstudy.com/updates/53f55faee4b0e0a0d4f63f1f Work by @IMStuck

Directrix (directrix):

Read it and let me know what you think, okay. @hellokitty01

OpenStudy (anonymous):

that makes sense. I did write a written explanation down but I don't think it's very good and I was trying to figure out the equation to use it as an example

OpenStudy (anonymous):

I guess I could come up with my own example to plug into the equation 1/11 + 1/14 + 1/m = 1/x for x and then show how Matthew's time could be found from that

Directrix (directrix):

Yes. You need to say what x represents. I think it is the time required if all three people worked together. Did you follow how to write the explanation for Matthew based on the other thread?

OpenStudy (anonymous):

yes i do

Directrix (directrix):

I don't know that you have to pick a number for Matthew to show him anything. It seems that if he knew he would need to solve for m, he might follow that. If he is like me, Matthew would wonder where the equation came from to start with.

OpenStudy (anonymous):

that's true. I'll just stick with the explanation then. thank you again!

Directrix (directrix):

If you want to post it here, I'll read it tomorrow if you like.

OpenStudy (anonymous):

ok :)